6.       You are considering a hotel purchase. The current purchase price is $ 3,000,000. The bank is willing to finance 70% of the purchase price for 20 years in quarterly installments at 8% per annum mortgage nominal rate. What is the balloon payment you will have to pay, if you want to resell the hotel after 10 years? Consider a fully amortizing fixed rate mortgage.

a)      $ 0

b)      $ 1,445,395

c)       $ 1,669,270

d)      $ 1,981,333
Answer #1

Step 1: Monthly payment
PV=-3000000*70%
N=20*4
I/Y=8%/4
FV=0
CPT PMT=52837.4814591

Step 2: Loan outstanding after 10 years
PMT=52837.4814591
PV=-3000000*70%
N=10*4
I/Y=8%/4
CPT FV=1445394.6271858

Balloon payment=Loan outstanding after 10 years=1445394.6271858